If the perimeter of a rectangle is 42 cm and the area is 90 square cm, then what are the dimensions of the rectangle?
Answer by Alan3354(69443)

If the perimeter of a rectangle is 42 cm and the area is 90 square cm, then what are the dimensions of the rectangle?
----------------
P = 2L + 2W = 42
L + W = 21
L*W = 90
---
Find a pair of factors of 90 with a sum of 21.